OpenStudy (vortish):

Find the point slope of the line having the given slope and containing the given point m = -5;(2,1)

OpenStudy (anonymous):

use this formula y-y1=m(x-x1) y1=1 x1=2

OpenStudy (anonymous):

u got it?

OpenStudy (vortish):

no im still having trouble with it

OpenStudy (anonymous):